Four protons form a square of side length a . what is the electric potential in the center of the square ?
1
Expert's answer
2020-06-29T14:04:37-0400

Let us denote the charge on a proton is "p=1.602\times 10^{-19}C"

Now, each sides of the square is "a" ,thus distance between each corner from the center of the square will be

"\frac{a}{\sqrt{2}}"

Potential due to each charge is

"V_i=\frac{1}{4\pi\epsilon_0}\frac{p}{\frac{a}{\sqrt{2}}}=\frac{1}{4\pi\epsilon_0}\frac{\sqrt{2}p}{a}"

Thus, potential at center of the square is

"\sum_{i=1}^{4}V_i=\frac{\sqrt{2}}{\pi\epsilon_0}\frac{p}{a}"